Help to solve this equation:
〖log〗_10 x+ 〖log〗_10 2= 〖log〗_10 10-〖log〗_10 1

Answer by stanbon(75887)

〖log〗_10 x+ 〖log〗_10 2= 〖log〗_10 10-〖log〗_10 1
-----
log(x) - log(2) = log(10) - log(1)
------
log(x/2) = 1
-------
x/2 = 10
x = 20
===============
Cheers,
Stan H.
---------------